OpenStudy (anonymous):

Express as a simplified polynomial: (4a + 5)^2

OpenStudy (anonymous):

\[(4a+5)^2=(4a+5)(4a+5)=16a^2+20a+20a+25\] \[=16a^2+40a+25\]

OpenStudy (anonymous):

or you can use \[(a+b)^2=a^2+2ab+b^2\] with \[a=4a,b=5\] it amounts to the same thing